# Computational Modeling for Personalized Transcranial Electrical Stimulation: Theory, Tools, and Applications

### v1

@@ 6 Conclusion
The tES is an effective non-invasive neuromodulation technique for treating neurological diseases and enhancing cognitive performance. Personalized tES implementation is crucial for mitigating variations in electric field distribution resulting from anatomical differences. Forward modeling and inverse optimization are vital stages in the computational modeling of personalized tES, facilitating the refinement of the stimulation strategy. Optimization methods can be utilized to design stimulation strategies that are more focused or multi-targeted. However, how the external tES interacts with the internal brain’s activity remains a topic of ongoing discussion. Combining tES with neural signal recording holds promise for analyzing stimulation results and designing more interpretative models.
@@ Acknowledgement
I would like to express our sincere gratitude to Dr. Zhichao Liang and Dr. Xinke Shen for their invaluable suggestions during the preparation of this manuscript.

### v2

@@ 6 Conclusion
tES is an effective non-invasive neuromodulation technique for treating neurological diseases and enhancing cognitive performance. Personalized tES implementation is crucial for mitigating variations in electric field distribution resulting from anatomical differences. Forward modeling and inverse optimization are vital stages in the computational modeling of personalized tES, facilitating the refinement of the stimulation strategy. Optimization methods can be utilized to design stimulation strategies that are more focused or multi-targeted. Integrating personalized tES computational modeling with neuroimaging and neurodynamic modeling offers excellent potential for enhancing stimulation precision. However, how the external tES interacts with the internal brain’s activity remains a topic of ongoing discussion. Combining tES with multimodal, high-precision personalized computational models is expected to become a key direction for future research.
@@ Acknowledgement
We would like to express our sincere gratitude to Dr. Zhichao Liang and Dr. Xinke Shen for their invaluable suggestions during the preparation of this manuscript. This work was supported by the National Natural Science Foundation of China (62472206), National Key R&D Program of China (2025YFC3410000), Shenzhen Science and Technology Innovation Committee (RCYX20231211090405003, KJZD20230923115221044), Guangdong Provincial Key Laboratory of Advanced Biomaterials (2022B1212010003), and the open research fund of the Guangdong Provincial Key Laboratory of Mathematical and Neural Dynamical Systems, the Center for Computational Science and Engineering at Southern University of Science and Technology.